### v1

# A Quantum Perspective on Uniqueness Pairs for the Fourier Transform

Abstract.
Kulikov, Nazarov, and Sodin recently introduced the notion of criticality to analyze when discrete subsets of the real line form uniqueness pairs for the Fourier transform, relying crucially on estimates derived from the Wirtinger–Poincaré inequality. In this work, drawing analogies from quantum mechanics, we propose an alternative approach to studying uniqueness sets in a Fourier-symmetric Sobolev space. Our method is based on eigenvalue estimates and offers a simplified perspective on the problem.
The Fourier-symmetric Sobolev space
$\mathcal{H}=\left\{f\>\Big|\>\int_{\mathbb{R}}(1+x^{2})|f(x)|^{2}\textnormal{d}x+\int_{\mathbb{R}}(1+k^{2})|\hat{f}(k)|^{2}\textnormal{d}k<\infty\right\}$
has recently attracted attention in the study of crystalline measures and sampling theory. In this note, we examine this space from a quantum-mechanical perspective, recovering recent results about uniqueness sets associated with the Fourier transform. We employ the “unitary convention” for the Fourier transform
$\hat{f}(k)=\mathcal{F}[f](k)=\frac{1}{\sqrt{2\pi}}\int_{\mathbb{R}}f(x)e^{-ikx}\textnormal{d}x,\quad f(x)=\mathcal{F}^{-1}[\hat{f}](x)=\frac{1}{\sqrt{2\pi}}\int_{\mathbb{R}}\hat{f}(k)e^{ikx}\textnormal{d}k.$
It will be convenient to view $\mathcal{H}$ as a subspace of $L^{2}(\mathbb{R})$, by introducing the norm
[…]
$=\int_{\mathbb{R}}x^{2}|f(x)|^{2}\textnormal{d}x+\int_{\mathbb{R}}|f^{\prime}(x)|^{2}\textnormal{d}x$
$=\int_{\mathbb{R}}\overline{f(x)}x^{2}f(x)\textnormal{d}x-\int_{\mathbb{R}}\overline{f(x)}\frac{\partial^{2}f}{\partial x^{2}}(x)dx$
$=2\int_{\mathbb{R}}\bar{f}H(f)\textnormal{d}x,$
where
$H=\frac{1}{2}\left(x^{2}-\frac{\partial^{2}}{\partial x^{2}}\right).$
$H$ is the Hamiltonian operator of the classical quantum harmonic oscillator [S]. The eigenvalue-eigenvector equation $Hf=Ef$ is the Schrödinger equation of the system [S]. Our computation assumes that the second derivative of $f$ exists. However, for the construction of $\mathcal{H}$ this is immaterial.
Theorem 1.
The following assertions hold for the Fourier-symmetric Sobolev space $\mathcal{H}$:
  1. (1)
    $\mathcal{H}\subset L^{2}(\mathbb{R})$ and $\mathcal{H}$ is the domain and image of the self-adjoint and positive operator $H=\frac{1}{2}\left(x^{2}-\frac{\partial^{2}}{\partial x^{2}}\right)$.
  2. (2)
    $H$ has eigenvectors that are Hermite functions of the form
    $\phi_{n}(x)=\frac{1}{\sqrt{n!}}\left(\frac{1}{\sqrt{2}}(x-\partial_{x})\right)^{n}\phi_{0},\ n\geq 0,$
    with $\phi_{0}=\pi^{-1/4}e^{-x^{2}/2}$. The respective eigenvalues are $E_{n}=(n+\frac{1}{2})$.
  3. (3)
    $H$ is self-dual under the Fourier transform: $\hat{H}=\mathcal{F}H\mathcal{F}^{-1}=H$.
  4. (4)
    The Hermite functions form an orthogonal basis for $\mathcal{H}$.
  5. (5)
[…]
and the fact that the eigenvalues of $H$ carry no degeneracy.
A central theme of our analysis is the spectral behavior of Hamiltonians confined by infinite potential barriers. We write for the Hamiltonian of the confined system
$H_{[a,b]}=\begin{cases}H\ \textnormal{if}\ x\in[a,b],\\ \infty\ \textnormal{else},\end{cases}$
in a distributional sense. This corresponds to the Schrödinger equation with Dirichlet boundary conditions
(0.1) $Hf=Ef\ \textnormal{under}\ f(a)=f(b)=0.$
Similarly to the unconfined harmonic oscillator, the solutions of the confined system (0.1) have been studied exhaustively, e.g. [BS]. With the boundary conditions $H_{[a,b]}$ is a self-adjoint and positive definite operator on $L^{2}{[a,b]}$. We will write $E_{[a,b]}^{(n)}$ for the eigenvalues of the confined system and $\psi_{[a,b]}^{(n)}$ for the respective eigenvectors. It holds automatically that $E_{[a,b]}^{(n)}>E_{n}$ because the introduction of confinement can only lead to an increase of energy levels.
In the context of uniqueness sets of the Fourier transform and crystalline measures the concept of criticality has recently emerged as a means to characterize the supports of these measures. The following definition orients itself at [KNS]. Item (1) is taken from there [^1] The notions of super- and subcriticality studied in [KNS] are more general in that we restrict ourselves to the symmetric situation $(p,q)=(2,2)$, cf. Def. 2 in [KNS]. This is to achieve a lean presentation., whereas (2) is original to this note.
Definition 2.
Let $(\Lambda,M)$ be a pair of countable subsets of $\mathbb{R}$. Enumerate the elements of $\Lambda$ and $M$ in increasing order.
  1. (1)
    The pair $(\Lambda,M)$ is uniformly supercritical if
    $\sup_{j\in\mathbb{Z}}\max\{|\lambda_{j}|,|\lambda_{j+1}|\}(\lambda_{j+1}-\lambda_{j})<\pi\ \textnormal{and}\ \sup_{j\in\mathbb{Z}}\max\{|\mu_{j}|,|\mu_{j+1}|\}(\mu_{j+1}-\mu_{j})<\pi.$
    The pair $(\Lambda,M)$ is supercritical if
    $\lim\sup|\lambda_{j}|(\lambda_{j+1}-\lambda_{j})<\pi\ \textnormal{and}\ \lim\sup|\mu_{j}|(\mu_{j+1}-\mu_{j})<\pi.$
    The pair $(\Lambda,M)$ is subcritical if
    $\lim\inf|\lambda_{j}|(\lambda_{j+1}-\lambda_{j})>\pi\ \textnormal{and}\ \lim\inf{|\mu_{j}|}(\mu_{j+1}-\mu_{j})>\pi.$
  2. (2)
    The pair $(\Lambda,M)$ has uniformly supercritical ground state energy if
    $\inf_{j\in\mathbb{Z}}\frac{\sqrt{E_{[\lambda_{j},\lambda_{j+1}]}^{(0)}}}{\max\{|\lambda_{j}|,|\lambda_{j+1}|\}}>1\ \textnormal{and}\ \inf_{j\in\mathbb{Z}}\frac{\sqrt{E_{[\mu_{j},\mu_{j+1}]}^{(0)}}}{\max\{|\mu_{j}|,|\mu_{j+1}|\}}>1,$
    where $E_{[a,b]}^{(0)}$ stands for the ground state energy of the Schrödinger equation (0.1). We define the notions of super- and subcritical ground state energy, mutatis mutandis, as in item (1).
Remark 3.
[…]
$\sup_{j\in\mathbb{Z}}\max\{|\tilde{\lambda}_{j}|,|\tilde{\lambda}_{j+1}|\}(\tilde{\lambda}_{j+1}-\tilde{\lambda}_{j})<\frac{1}{2}\ \textnormal{and}\ \sup_{j\in\mathbb{Z}}\max\{|\tilde{\mu}_{j}|,|\tilde{\mu}_{j+1}|\}(\tilde{\mu}_{j+1}-\tilde{\mu}_{j})<\frac{1}{2}.$
This discrepancy is due to the choice of Fourier convention. To transcript between the conventions consider $M^{\prime}=2\pi\tilde{M}$ such that $\hat{f}(\mu_{i}^{\prime})=0$ for all $\mu_{j}^{\prime}\in M^{\prime}$ under our convention. The respective criticality condition for $M^{\prime}$ reads $\sup_{j\in\mathbb{Z}}\max\{|\mu_{j}^{\prime}|,|\mu_{j+1}^{\prime}|\}(\mu_{j+1}^{\prime}-\tilde{\mu}_{j}^{\prime})<\frac{1}{2}\cdot(2\pi)^{2}$. $(\tilde{\Lambda},M^{\prime})$ is a uniqueness set if and only if $(\alpha\tilde{\Lambda},\frac{1}{\alpha}M^{\prime})$, $\alpha\neq 0$ is a uniqueness set. We use this property setting $\Lambda=\sqrt{2\pi}\tilde{\Lambda}$ and $M=\frac{1}{\sqrt{2\pi}}M^{\prime}$.
The following theorem takes a pivotal role in the development of the theory of criticality. It serves as a “toy model” to outline the main line of reasoning in a way that is “short and free of cumbersome technical details” in [KNS].
Theorem 4 (Theorem 1A [KNS]).
Let $(\Lambda,M)$ be a uniformly supercritical pair. Then the only function $f\in\mathcal{H}$ such that $f(\lambda_{i})=\hat{f}(\mu_{i})=0$ for all $i\in\mathbb{Z}$ is $f=0$.
An essential part in the proof of the theorem is played the Wirtinger–Poincaré inequality. We will use this theorem as a cornerstone to guide our further investigation. But rather than making use of the Wirtinger–Poincaré inequality, we will employ spectral methods to arrive at our estimates. Our definition of uniformly supercritical ground state energy is motivated by the following spectral estimate on the Rayleigh quotient of the confined Hamiltonian. It holds that
$\frac{\int_{a}^{b}\bar{f}H_{[a,b]}f\textnormal{d}x}{\int_{a}^{b}|f|^{2}\textnormal{d}x}\geq E_{[a,b]}^{(0)},\ f\neq 0.$
This leads to the following theorem that mimics Theorem 4.
Theorem 5.
Let the pair $(\Lambda,M)$ have uniformly supercritical ground energy. Then the only function $f\in\mathcal{H}$ such that $f(\lambda_{i})=\hat{f}(\mu_{i})=0$ for all $i\in\mathbb{Z}$ is $f=0$.
Proof.
We make use of Theorem 1, item (1), and of the unitarity of the Fourier transform. We phrase the confined Schrödinger equation as an eigenvalue problem of a confined Hamiltonian. We then use the mentioned lower bound on the Rayleigh quotient to estimate expectations of $H_{[a,b]}$ and $\hat{H}_{[a,b]}$. Recall that $H=\hat{H}$ according to Theorem 4, item (3). This yields
[…]
$=\int_{\mathbb{R}}\bar{f}Hf\textnormal{d}x+\int_{\mathbb{R}}\overline{\hat{f}}\hat{H}\hat{f}\textnormal{d}k$
$=\sum_{j\in\mathbb{Z}}\left(\int_{\lambda_{j}}^{\lambda_{j+1}}\bar{f}Hf\textnormal{d}x+\int_{\mu_{j}}^{\mu_{j+1}}\overline{\hat{f}}\hat{H}\hat{f}\textnormal{d}k\right)$
$=\sum_{j\in\mathbb{Z}}\left(\int_{\lambda_{j}}^{\lambda_{j+1}}{\bar{f}}H_{[\lambda_{j},\lambda_{j+1}]}f\textnormal{d}x+\int_{\mu_{j}}^{\mu_{j+1}}\overline{\hat{f}}\hat{H}_{[\mu_{j},\mu_{j+1}]}\hat{f}\textnormal{d}k\right)$
$\geq\sum_{j\in\mathbb{Z}}\left(E_{[\lambda_{j},\lambda_{j+1}]}^{(0)}\int_{\lambda_{j}}^{\lambda_{j+1}}|f|^{2}\textnormal{d}x+E_{[\mu_{j},\mu_{j+1}]}^{(0)}\int_{\mu_{j}}^{\mu_{j+1}}|\hat{f}|^{2}\textnormal{d}k\right)$
$\geq\sum_{j\in\mathbb{Z}}\left(\frac{E_{[\lambda_{j},\lambda_{j+1}]}^{(0)}}{\max\{|\lambda_{j}|^{2},|\lambda_{j+1}|^{2}\}}\int_{\lambda_{j}}^{\lambda_{j+1}}x^{2}|f|^{2}\textnormal{d}x+\frac{E_{[\mu_{j},\mu_{j+1}]}^{(0)}}{\max\{|\mu_{j}|^{2},|\mu_{j+1}|^{2}\}}\int_{\mu_{j}}^{\mu_{j+1}}k^{2}|\hat{f}|^{2}\textnormal{d}k\right).$
Akin to assuming uniformly supercritical ground state energy of $(\Lambda,M)$, we assume that the quantities in Definition 2, (4), are bounded from below by $l$. It follows that
[…]
If $l>1$ this is a contradiction unless $f=0$. ∎
We see three advantages in Theorem 5 over Theorem 4: 1) Obtaining lower estimates on the energy levels $E_{[a,b]}^{(n)}$ is straight forward. Their analytic form can be derived using perturbation theory. 2) Using eigenvalue bounds, the proof of Theorem 5 provides a direct route to sharpness, with the lower bound on the Rayleigh quotient attained by the ground-state eigenfunction. 3) The derivation of Theorem 5 is simpler than the one of Theorem 4 presented in [KNS].
We begin by studying the energy levels. Assuming $0\notin[a,b]$, the energy levels of the confined harmonic oscillators can be bounded by the respective energy levels of a particle in constant elevated box potential
$H_{[a,b]}^{down}=\begin{cases}-\frac{1}{2}\frac{\partial^{2}}{\partial x^{2}}+\frac{1}{2}\min\{a^{2},b^{2}\}\ \textnormal{if}\ x\in[a,b],\\ \infty\ \textnormal{else},\end{cases}\quad\textnormal{and}\quad H_{[a,b]}>H_{[a,b]}^{down},$
$H_{[a,b]}^{up}=\begin{cases}-\frac{1}{2}\frac{\partial^{2}}{\partial x^{2}}+\frac{1}{2}\max\{a^{2},b^{2}\}\ \textnormal{if}\ x\in[a,b],\\ \infty\ \textnormal{else},\end{cases}\quad\textnormal{and}\quad H_{[a,b]}<H_{[a,b]}^{up}.$
These systems have the following eigenvectors for $n\geq 0$
$\psi_{[a,b]}^{(n)}=\sqrt{\frac{2}{b-a}}\sin\left(\frac{\pi(n+1)(x-a)}{b-a}\right),$
with respective eigenvalues
$E_{[a,b]}^{down,(n)}=\frac{(n+1)^{2}\pi^{2}}{2(b-a)^{2}}+\frac{1}{2}\min\{a^{2},b^{2}\}\ \textnormal{and}\ E_{[a,b]}^{up,(n)}=\frac{(n+1)^{2}\pi^{2}}{2(b-a)^{2}}+\frac{1}{2}\max\{a^{2},b^{2}\}.$
For the ground state energy of the confined Harmonic oscillators this implies
$\frac{\pi^{2}}{2(b-a)^{2}}+\frac{1}{2}\max\{a^{2},b^{2}\}>E^{(0)}_{[a,b]}>\frac{\pi^{2}}{2(b-a)^{2}}+\frac{1}{2}\min\{a^{2},b^{2}\}.$
To compare theorems 4 and 5 we show that the notions of criticality and criticality of the ground state energy are closely related.
Lemma 6.
The following assertions hold for a pair $(\Lambda,M)$:
  1. (1)
    If $(\Lambda,M)$ has uniformly supercritical ground state energy then $(\Lambda,M)$ is uniformly supercritical.
  2. (2)
    If $(\Lambda,M)$ is uniformly supercritical, then $(\Lambda,M)$ has uniformly supercritical ground state energy with the exception of maximally a finite number of intervals $\{[\lambda_{i},\lambda_{i+1}]\}_{i}$, $\{[\mu_{j},\mu_{j+1}]\}_{i}$ with $|i|<N_{\Lambda}$ and $|j|<N_{M}$.
  3. (3)
    $\lim\inf\frac{\sqrt{E_{[\lambda_{j},\lambda_{j+1}]}^{(0)}}}{\max\{|\lambda_{j}|,|\lambda_{j+1}|\}}$ $=\frac{\pi}{\lim\sup(\lambda_{j+1}-\lambda_{j})\max\{|\lambda_{j}|,|\lambda_{j+1}|\}},$
    $\lim\inf\frac{\sqrt{E_{[\mu_{j},\mu_{j+1}]}^{(0)}}}{\max\{|\mu_{j}|,|\lambda_{j+1}|\}}$ $=\frac{\pi}{\lim\sup(\mu_{j+1}-\mu_{j})\max\{|\mu_{j}|,|\mu_{j+1}|\}}.$
Proof.
We show the assertions for $\Lambda$; the same reasoning applies to $M$.
  1. (1)
    Making use of our upper bound on the ground state energy and the condition $\frac{{E_{[\lambda_{j},\lambda_{j+1}]}^{(0)}}}{\max\{|\lambda_{j}|^{2},|\lambda_{j+1}|^{2}\}}>1$, we have
    $1<$ $\frac{{E_{[\lambda_{j},\lambda_{j+1}]}^{(0)}}}{\max\{|\lambda_{j}|^{2},|\lambda_{j+1}|^{2}\}}<\frac{\pi^{2}}{{2}(\lambda_{j+1}-\lambda_{j})^{2}\max\{|\lambda_{j}|^{2},|\lambda_{j+1}|^{2}\}}+\frac{1}{2},$
    which implies $\sup_{j\in\mathbb{Z}}\max\{|\lambda_{j}|,|\lambda_{j+1}|\}(\lambda_{j+1}-\lambda_{j})<\pi$.
  2. (2)
    Making use of our lower bound on the ground state energy and the condition $\max\{|\lambda_{j}|,|\lambda_{j+1}|\}(\lambda_{j+1}-\lambda_{j})<\pi$, we have
    $\frac{{E_{[\lambda_{j},\lambda_{j+1}]}^{(0)}}}{\max\{|\lambda_{j}|^{2},|\lambda_{j+1}|^{2}\}}\geq\frac{\pi^{2}}{{2}(\lambda_{j+1}-\lambda_{j})^{2}\max\{|\lambda_{j}|^{2},|\lambda_{j+1}|^{2}\}}+\frac{\min\{|\lambda_{j}|^{2},|\lambda_{j+1}|^{2}\}}{2\max\{|\lambda_{j}|^{2},|\lambda_{j+1}|^{2}\}}$
    $>\frac{1}{2}+\frac{1}{2}\frac{\min\{|\lambda_{j}|^{2},|\lambda_{j+1}|^{2}\}}{\max\{|\lambda_{j}|^{2},|\lambda_{j+1}|^{2}\}}.$
    It is obvious that if $|\lambda_{j+1}-\lambda_{j}|\rightarrow 0$ then the second summand approaches $\frac{1}{2}$ and consequently in this case $\inf_{j\in\mathbb{Z}}\frac{\sqrt{E_{[\lambda_{j},\lambda_{j+1}]}^{(0)}}}{\max\{|\lambda_{j}|,|\lambda_{j+1}|\}}>1$ with the possible exception of a finite number of terms. Hence, we show that $|\lambda_{j+1}-\lambda_{j}|\rightarrow 0$. Using standard $\theta(\cdot)$ notation from asymptotic analysis, suppose that $|\lambda_{n}|=\theta(|n|^{\alpha})$, $\alpha>0$. Then making use of a telescope sum we find
    $|\lambda_{n}-\lambda_{0}|\leq\sum_{i=0}^{n-1}|\lambda_{i+1}-\lambda_{i}|\leq\frac{1}{2}\sum_{i=0}^{N-1}\theta(|n|^{-\alpha}).$
    If $\alpha>1$ then the right hand side converges, while the left hand side grows, which is a contradiction. Hence $\alpha\leq 1$. It can be argued similarly that $\alpha<1$. In this case $\lambda_{n+1}-\lambda_{n}\rightarrow 0$ when $n\rightarrow\infty$.
  3. (3)
    This follows from the above proofs of (1) and (2).
In view of Theorem 5 and Lemma 6, the approach of [KNS] can now be revisited, and their results recovered within our framework. As this would amount to a repetition of existing work, we present this note primarily as a conceptual contribution.
Acknowledgements
This work was carried out during a research stay hosted by Andrii Bondarenko and Kristian Seip. The author is grateful to them for valuable discussions, and also wishes to thank the Department of Mathematical Sciences at NTNU for their warm hospitality and supportive environment. This work was supported by the Swiss National Science foundation, SNF grant No. CRSK-2_229036.