### v2

# Spectral Criteria for Unique Signal Recovery from Two-Sided Sampling

Abstract.
The identification of sampling sets that enable unique signal recovery is fundamental to many applications in signal processing and remains a central problem in mathematical analysis. Recent studies in the mathematical literature, particularly in the context of the Fourier transform and crystalline measures, have developed a theory that empowers signal recovery from two-sided sampling in both time and frequency domains. Kulikov, Nazarov, and Sodin introduced a method for identifying pairs of sets that enable unique recovery, based on functional inequalities of the Wirtinger-Poincaré type. In this work, we propose an alternative, spectral approach based on analogies with quantum mechanics. By relating uniqueness pairs to eigenvalue estimates of associated self-adjoint operators, our method offers a conceptually simpler and more flexible framework for studying signal recovery from two-sided sampling. Our approach extends naturally to other unitary transforms commonly used in signal processing. We demonstrate its effectiveness in the contexts of the Fractional Fourier transform and the Hankel transform.
The ability to reconstruct a function from its values on a discrete set lies at the heart of sampling theory, a cornerstone of both digital signal processing and applied mathematics. A fundamental application of this theory is the conversion between digital and analog signals. A central role in sampling theory is played by the concept of a uniqueness set, which generalizes the classical Nyquist rate of Whittaker-Kotelnikov-Shannon sampling theory beyond Paley-Wiener spaces. The characterization of uniqueness sets is closely tied to the geometry of the sampling space and the density of sampling points [S], making their study foundational to both theory and applications.
Recently, in the mathematical literature, particularly within the context of the Fourier transform and crystalline measures, the concept of a uniqueness pair has emerged. This pair consists of two sets: one in the real domain and the other in the Fourier domain, with the property that if values are specified for an interpolating function at these sets, this function is determined uniquely, see [RV], and also [RS, KNS]. Beyond its mathematical significance, this concept has practical implications in engineering, as it enables the unique recovery of a signal from two-sided sampling - i.e., when samples are concurrently taken in both the real and Fourier domains. Achieving this typically requires that the sampling densities in both domains surpass certain critical thresholds, the exact form of which depends on the geometry of the underlying spaces.
Current approaches to determining these recovery thresholds rely crucially on estimates derived from the Wirtinger–Poincaré inequality [KNS]. Motivated by analogies with quantum mechanics, we propose an alternative spectral approach to the analysis of uniqueness pairs. Our framework offers a natural and flexible method for addressing uniqueness problems for the Fourier transform and extends directly to a broad class of unitary transformations on the real line.
Beyond the classical Fourier transform, several related unitary transforms play a fundamental role in signal processing and engineering applications. We extend the idea of two-sided sampling to these settings and we provide a methodology to study uniqueness pairs for such transformations. To illustrate the effectiveness of our approach, we consider the Fractional Fourier and the Hankel transforms. The Fractional Fourier transform provides a continuous interpolation between the time and frequency domains and is widely used in optical signal analysis, radar imaging, and time-frequency filtering [OZK]. The Hankel transform arises naturally in problems with radial symmetry, including wave propagation, diffraction, and medical imaging such as computer tomography [KS].
1. Uniqueness pairs for the Fourier transform
We begin by considering the Fourier-symmetric Sobolev space
$\mathcal{H}=\left\{f\>\Big|\>\int_{\mathbb{R}}(1+x^{2})|f(x)|^{2}\textnormal{d}x+\int_{\mathbb{R}}(1+k^{2})|\hat{f}(k)|^{2}\textnormal{d}k<\infty\right\},$
which arises naturally in the theory of modulation spaces - function spaces used in time-frequency analysis to quantify the joint localization of a signal in both time and frequency; see, e.g., [EG]. More recently, $\mathcal{H}$ has also attracted attention in the study of crystalline measures and two-sided sampling theory [KNS, Z]. In this section, we examine $\mathcal{H}$ from a quantum-mechanical perspective and recover several recent results concerning uniqueness pairs associated with the Fourier transform. We adopt the unitary convention for the Fourier transform, as it provides conceptual clarity and facilitates the generalization of the notion of uniqueness pairs to other unitary transformations:
$\hat{f}(k)={F}[f](k)=\frac{1}{\sqrt{2\pi}}\int_{\mathbb{R}}f(x)e^{-ikx}\textnormal{d}x,\quad f(x)={F}^{-1}[\hat{f}](x)=\frac{1}{\sqrt{2\pi}}\int_{\mathbb{R}}\hat{f}(k)e^{ikx}\textnormal{d}k.$
We note that the convention used in [KNS] is not unitary. It will be convenient to view $\mathcal{H}$ as a subspace of $L^{2}(\mathbb{R})$, by introducing the norm
[…]
$=\int_{\mathbb{R}}x^{2}|f(x)|^{2}\textnormal{d}x+\int_{\mathbb{R}}|f^{\prime}(x)|^{2}\textnormal{d}x$
$=\int_{\mathbb{R}}\overline{f(x)}x^{2}f(x)\textnormal{d}x-\int_{\mathbb{R}}\overline{f(x)}\frac{\partial^{2}f}{\partial x^{2}}(x)dx$
$=2\int_{\mathbb{R}}\bar{f}H(f)\textnormal{d}x,$
where
$H=\frac{1}{2}\left(x^{2}-\frac{\partial^{2}}{\partial x^{2}}\right).$
$H$ is the Hamiltonian operator of the classical quantum harmonic oscillator [NS]. The eigenvalue-eigenvector equation $Hf=Ef$ is the Schrödinger equation of the system [NS]. Our computation assumes that the second derivative of $f$ exists. However, for the construction of $\mathcal{H}$ this is immaterial.
Theorem 1.
The following assertions hold for the Fourier-symmetric Sobolev space $\mathcal{H}$:
  1. (1)
    $\mathcal{H}\subset L^{2}(\mathbb{R})$ and $\mathcal{H}$ is the domain and image of an essentially self-adjoint and positive operator $H=\frac{1}{2}\left(x^{2}-\frac{\partial^{2}}{\partial x^{2}}\right)$.
  2. (2)
    $H$ has eigenvectors that are Hermite functions of the form
    $\phi_{n}(x)=\frac{1}{\sqrt{n!}}\left(\frac{1}{\sqrt{2}}(x-\partial_{x})\right)^{n}\phi_{0}(x),\ n\geq 0,$
    with $\phi_{0}=\pi^{-1/4}e^{-x^{2}/2}$. The respective eigenvalues are $E_{n}=(n+\frac{1}{2})$.
  3. (3)
    $H$ is self-dual under the Fourier transform: $\hat{H}={F}H{F}^{-1}=H$.
  4. (4)
    The Hermite functions form an orthogonal basis for $\mathcal{H}$.
  5. (5)
[…]
and the fact that the eigenvalues of $H$ carry no degeneracy.
A central theme of our analysis is the spectral behavior of Hamiltonians confined by infinite potential barriers. We write for the Hamiltonian of the confined system
$H_{[a,b]}=\begin{cases}H\ \textnormal{if}\ x\in[a,b],\\ \infty\ \textnormal{else},\end{cases}$
in a distributional sense. This corresponds to the Schrödinger equation with Dirichlet boundary conditions
(1.1) $Hf=Ef\ \textnormal{under}\ f(a)=f(b)=0.$
Similarly to the unconfined harmonic oscillator, the solutions of the confined system (1.1) have been studied exhaustively, e.g. [BS]. With the boundary conditions $H_{[a,b]}$ is a self-adjoint and positive definite operator on $L^{2}{[a,b]}$. We will write $E_{[a,b]}^{(n)}$ for the eigenvalues of the confined system and $\psi_{[a,b]}^{(n)}$ for the respective eigenvectors. It holds automatically that $E_{[a,b]}^{(n)}>E_{n}$ because the introduction of confinement can only lead to an increase of energy levels. The following definition follows [KNS]. Item (1) is taken from there [^1] The notions of super- and subcriticality studied in [KNS] are more general in that we restrict ourselves to the symmetric situation $(p,q)=(2,2)$, cf. Def. 2 in [KNS]. This is to achieve a lean presentation., whereas (2) is original to this note.
Definition 2.
Let $(\Lambda,M)$ be a pair of countable subsets of $\mathbb{R}$. Enumerate the elements of $\Lambda$ and $M$ in increasing order, denoting them by $(\lambda_{j})_{j\in\mathbb{Z}}$ and $(\mu_{j})_{j\in\mathbb{Z}}$, respectively.
  1. (1)
    The pair $(\Lambda,M)$ is uniformly supercritical if
    $\sup_{j\in\mathbb{Z}}\max\{|\lambda_{j}|,|\lambda_{j+1}|\}(\lambda_{j+1}-\lambda_{j})<\pi\ \textnormal{and}\ \sup_{j\in\mathbb{Z}}\max\{|\mu_{j}|,|\mu_{j+1}|\}(\mu_{j+1}-\mu_{j})<\pi.$
    The pair $(\Lambda,M)$ is supercritical if
    $\lim\sup|\lambda_{j}|(\lambda_{j+1}-\lambda_{j})<\pi\ \textnormal{and}\ \lim\sup|\mu_{j}|(\mu_{j+1}-\mu_{j})<\pi.$
    The pair $(\Lambda,M)$ is subcritical if
    $\lim\inf|\lambda_{j}|(\lambda_{j+1}-\lambda_{j})>\pi\ \textnormal{and}\ \lim\inf{|\mu_{j}|}(\mu_{j+1}-\mu_{j})>\pi.$
  2. (2)
    The pair $(\Lambda,M)$ has uniformly supercritical ground state energy if
    $\inf_{j\in\mathbb{Z}}\frac{\sqrt{E_{[\lambda_{j},\lambda_{j+1}]}^{(0)}}}{\max\{|\lambda_{j}|,|\lambda_{j+1}|\}}>1\ \textnormal{and}\ \inf_{j\in\mathbb{Z}}\frac{\sqrt{E_{[\mu_{j},\mu_{j+1}]}^{(0)}}}{\max\{|\mu_{j}|,|\mu_{j+1}|\}}>1,$
    where $E_{[a,b]}^{(0)}$ stands for the ground state energy of the Schrödinger equation (1.1). We define the notions of super- and subcritical ground state energy, mutatis mutandis, as in item (1).
Remark 3.
[…]
$\sup_{j\in\mathbb{Z}}\max\{|\tilde{\lambda}_{j}|,|\tilde{\lambda}_{j+1}|\}(\tilde{\lambda}_{j+1}-\tilde{\lambda}_{j})<\frac{1}{2}\ \textnormal{and}\ \sup_{j\in\mathbb{Z}}\max\{|\tilde{\mu}_{j}|,|\tilde{\mu}_{j+1}|\}(\tilde{\mu}_{j+1}-\tilde{\mu}_{j})<\frac{1}{2}.$
This discrepancy is due to the choice of Fourier convention. To transcript between the conventions consider $M^{\prime}=2\pi\tilde{M}$ such that $\hat{f}(\mu_{i}^{\prime})=0$ for all $\mu_{j}^{\prime}\in M^{\prime}$ under our convention. The respective criticality condition for $M^{\prime}$ reads $\sup_{j\in\mathbb{Z}}\max\{|\mu_{j}^{\prime}|,|\mu_{j+1}^{\prime}|\}(\mu_{j+1}^{\prime}-\tilde{\mu}_{j}^{\prime})<\frac{1}{2}\cdot(2\pi)^{2}$. $(\tilde{\Lambda},M^{\prime})$ is a uniqueness set if and only if $(\alpha\tilde{\Lambda},\frac{1}{\alpha}M^{\prime})$, $\alpha\neq 0$ is a uniqueness set. We use this property setting $\Lambda=\sqrt{2\pi}\tilde{\Lambda}$ and $M=\frac{1}{\sqrt{2\pi}}M^{\prime}$.
The following theorem takes a pivotal role in the development of the theory of criticality. It serves as a “toy model” to outline the main line of reasoning in a way that is “short and free of cumbersome technical details” in [KNS].
Theorem 4 (Theorem 1A [KNS]).
Let $(\Lambda,M)$ be a uniformly supercritical pair. Then the only function $f\in\mathcal{H}$ such that $f(\lambda_{i})=\hat{f}(\mu_{i})=0$ for all $i\in\mathbb{Z}$ is $f=0$.
The elegant proof presented in [KNS] relies crucially on the Wirtinger–Poincaré inequality. A closely related argument, involving the same inequality, has appeared in the context of irregular sampling in [G]. We will use this theorem as a cornerstone to guide our further investigation. But rather than making use of the Wirtinger–Poincaré inequality, we will employ spectral methods to arrive at our estimates. Our definition of uniformly supercritical ground state energy is motivated by the following spectral estimate on the Rayleigh quotient of the confined Hamiltonian. It holds that
$\frac{\int_{a}^{b}\bar{f}H_{[a,b]}f\textnormal{d}x}{\int_{a}^{b}|f|^{2}\textnormal{d}x}\geq E_{[a,b]}^{(0)},\ f\neq 0.$
This leads to the following theorem that mimics Theorem 4.
Theorem 5.
Let the pair $(\Lambda,M)$ have uniformly supercritical ground state energy. Then the only function $f\in\mathcal{H}$ such that $f(\lambda_{i})=\hat{f}(\mu_{i})=0$ for all $i\in\mathbb{Z}$ is $f=0$.
Proof.
We make use of Theorem 1, item (1), and of the unitarity of the Fourier transform. We phrase the confined Schrödinger equation as an eigenvalue problem of a confined Hamiltonian. We then use the mentioned lower bound on the Rayleigh quotient to estimate expectations of $H_{[a,b]}$ and $\hat{H}_{[a,b]}$. Recall that $H=\hat{H}$ according to Theorem 4, item (3). This yields
[…]
$=\int_{\mathbb{R}}\bar{f}Hf\textnormal{d}x+\int_{\mathbb{R}}\overline{\hat{f}}\hat{H}\hat{f}\textnormal{d}k$
$=\sum_{j\in\mathbb{Z}}\left(\int_{\lambda_{j}}^{\lambda_{j+1}}\bar{f}Hf\textnormal{d}x+\int_{\mu_{j}}^{\mu_{j+1}}\overline{\hat{f}}\hat{H}\hat{f}\textnormal{d}k\right)$
$=\sum_{j\in\mathbb{Z}}\left(\int_{\lambda_{j}}^{\lambda_{j+1}}{\bar{f}}H_{[\lambda_{j},\lambda_{j+1}]}f\textnormal{d}x+\int_{\mu_{j}}^{\mu_{j+1}}\overline{\hat{f}}\hat{H}_{[\mu_{j},\mu_{j+1}]}\hat{f}\textnormal{d}k\right)$
$\geq\sum_{j\in\mathbb{Z}}\left(E_{[\lambda_{j},\lambda_{j+1}]}^{(0)}\int_{\lambda_{j}}^{\lambda_{j+1}}|f|^{2}\textnormal{d}x+E_{[\mu_{j},\mu_{j+1}]}^{(0)}\int_{\mu_{j}}^{\mu_{j+1}}|\hat{f}|^{2}\textnormal{d}k\right)$
$\geq\sum_{j\in\mathbb{Z}}\left(\frac{E_{[\lambda_{j},\lambda_{j+1}]}^{(0)}}{\max\{|\lambda_{j}|^{2},|\lambda_{j+1}|^{2}\}}\int_{\lambda_{j}}^{\lambda_{j+1}}x^{2}|f|^{2}\textnormal{d}x+\frac{E_{[\mu_{j},\mu_{j+1}]}^{(0)}}{\max\{|\mu_{j}|^{2},|\mu_{j+1}|^{2}\}}\int_{\mu_{j}}^{\mu_{j+1}}k^{2}|\hat{f}|^{2}\textnormal{d}k\right).$
Akin to assuming uniformly supercritical ground state energy of $(\Lambda,M)$, we assume that the quantities in Definition 2, (4), are bounded from below by $l$. It follows that
[…]
If $l>1$ this is a contradiction unless $f=0$. ∎
We see three advantages in Theorem 5 over Theorem 4: 1) Lower bounds on the energy levels $E_{[a,b]}^{(n)}$ are straightforward to obtain. Their analytic form can be derived using perturbation theory. 2) The proof of Theorem 5, via eigenvalue bounds, gives a direct route to establishing sharpness, with the lower bound on the Rayleigh quotient attained by the ground-state eigenfunction. 3) The derivation of Theorem 5 is simpler than that of Theorem 4 as presented in [KNS], and it naturally extends to unitary transformations beyond the Fourier transform.
We begin by studying the energy levels. Assuming $0\notin[a,b]$, the energy levels of the confined harmonic oscillator can be bounded by the respective energy levels of a particle in constant elevated box potential
$H_{[a,b]}^{down}=\begin{cases}-\frac{1}{2}\frac{\partial^{2}}{\partial x^{2}}+\frac{1}{2}\min\{a^{2},b^{2}\}\ \textnormal{if}\ x\in[a,b],\\ \infty\ \textnormal{else},\end{cases}\quad\textnormal{and}\quad H_{[a,b]}>H_{[a,b]}^{down},$
$H_{[a,b]}^{up}=\begin{cases}-\frac{1}{2}\frac{\partial^{2}}{\partial x^{2}}+\frac{1}{2}\max\{a^{2},b^{2}\}\ \textnormal{if}\ x\in[a,b],\\ \infty\ \textnormal{else},\end{cases}\quad\textnormal{and}\quad H_{[a,b]}<H_{[a,b]}^{up}.$
These systems have the following eigenvectors for $n\geq 0$
$\psi_{[a,b]}^{(n)}=\sqrt{\frac{2}{b-a}}\exp\left(i\frac{\pi(n+1)(x-a)}{b-a}\right),$
with respective eigenvalues
$E_{[a,b]}^{down,(n)}=\frac{(n+1)^{2}\pi^{2}}{2(b-a)^{2}}+\frac{1}{2}\min\{a^{2},b^{2}\}\ \textnormal{and}\ E_{[a,b]}^{up,(n)}=\frac{(n+1)^{2}\pi^{2}}{2(b-a)^{2}}+\frac{1}{2}\max\{a^{2},b^{2}\}.$
For the ground state energy of the confined harmonic oscillator this implies
$\frac{\pi^{2}}{2(b-a)^{2}}+\frac{1}{2}\max\{a^{2},b^{2}\}>E^{(0)}_{[a,b]}>\frac{\pi^{2}}{2(b-a)^{2}}+\frac{1}{2}\min\{a^{2},b^{2}\}.$
To compare theorems 4 and 5 we show that the notions of criticality and criticality of the ground state energy are closely related.
Lemma 6.
The following assertions hold for a pair $(\Lambda,M)$:
  1. (1)
    If $(\Lambda,M)$ has uniformly supercritical ground state energy then $(\Lambda,M)$ is uniformly supercritical.
  2. (2)
    If $(\Lambda,M)$ is uniformly supercritical, then $(\Lambda,M)$ has uniformly supercritical ground state energy with the exception of maximally a finite number of intervals $\{[\lambda_{i},\lambda_{i+1}]\}_{i}$, $\{[\mu_{j},\mu_{j+1}]\}_{i}$ with $|i|<N_{\Lambda}$ and $|j|<N_{M}$.
  3. (3)
    The notions of supercriticality and supercriticality of the ground state energy are equivalent
    $\lim\inf\frac{\sqrt{E_{[\lambda_{j},\lambda_{j+1}]}^{(0)}}}{\max\{|\lambda_{j}|,|\lambda_{j+1}|\}}$ $=\frac{\pi}{\lim\sup(\lambda_{j+1}-\lambda_{j})\max\{|\lambda_{j}|,|\lambda_{j+1}|\}},$
    $\lim\inf\frac{\sqrt{E_{[\mu_{j},\mu_{j+1}]}^{(0)}}}{\max\{|\mu_{j}|,|\mu_{j+1}|\}}$ $=\frac{\pi}{\lim\sup(\mu_{j+1}-\mu_{j})\max\{|\mu_{j}|,|\mu_{j+1}|\}}.$
    The analogous statement holds also for subcriticality and supercriticality of the ground state energy.
Proof.
We show the assertions for $\Lambda$; the same reasoning applies to $M$.
  1. (1)
    Making use of our upper bound on the ground state energy and the condition $\frac{{E_{[\lambda_{j},\lambda_{j+1}]}^{(0)}}}{\max\{|\lambda_{j}|^{2},|\lambda_{j+1}|^{2}\}}>1$, we have
    $1<$ $\frac{{E_{[\lambda_{j},\lambda_{j+1}]}^{(0)}}}{\max\{|\lambda_{j}|^{2},|\lambda_{j+1}|^{2}\}}<\frac{\pi^{2}}{{2}(\lambda_{j+1}-\lambda_{j})^{2}\max\{|\lambda_{j}|^{2},|\lambda_{j+1}|^{2}\}}+\frac{1}{2},$
    which implies $\sup_{j\in\mathbb{Z}}\max\{|\lambda_{j}|,|\lambda_{j+1}|\}(\lambda_{j+1}-\lambda_{j})<\pi$.
  2. (2)
    Making use of our lower bound on the ground state energy and the condition $\max\{|\lambda_{j}|,|\lambda_{j+1}|\}(\lambda_{j+1}-\lambda_{j})<\pi$, we have
    $\frac{{E_{[\lambda_{j},\lambda_{j+1}]}^{(0)}}}{\max\{|\lambda_{j}|^{2},|\lambda_{j+1}|^{2}\}}\geq\frac{\pi^{2}}{{2}(\lambda_{j+1}-\lambda_{j})^{2}\max\{|\lambda_{j}|^{2},|\lambda_{j+1}|^{2}\}}+\frac{\min\{|\lambda_{j}|^{2},|\lambda_{j+1}|^{2}\}}{2\max\{|\lambda_{j}|^{2},|\lambda_{j+1}|^{2}\}}$
    $>\frac{1}{2}+\frac{1}{2}\frac{\min\{|\lambda_{j}|^{2},|\lambda_{j+1}|^{2}\}}{\max\{|\lambda_{j}|^{2},|\lambda_{j+1}|^{2}\}}.$
    It is obvious that if $|\lambda_{j+1}-\lambda_{j}|\rightarrow 0$ then the second summand approaches $\frac{1}{2}$ and consequently in this case $\inf_{j\in\mathbb{Z}}\frac{\sqrt{E_{[\lambda_{j},\lambda_{j+1}]}^{(0)}}}{\max\{|\lambda_{j}|,|\lambda_{j+1}|\}}>1$ with the possible exception of a finite number of terms. Hence, we show that $|\lambda_{j+1}-\lambda_{j}|\rightarrow 0$. Using standard $\theta(\cdot)$ notation from asymptotic analysis, suppose that $|\lambda_{n}|=\theta(|n|^{\alpha})$, $\alpha>0$. Then making use of a telescope sum we find
    $|\lambda_{n}-\lambda_{0}|\leq\sum_{i=0}^{n-1}|\lambda_{i+1}-\lambda_{i}|\leq\frac{1}{2}\sum_{i=0}^{N-1}\theta(|n|^{-\alpha}).$
    If $\alpha>1$ then the right hand side converges, while the left hand side grows, which is a contradiction. Hence $\alpha\leq 1$. It can be argued similarly that $\alpha<1$. In this case $\lambda_{n+1}-\lambda_{n}\rightarrow 0$ when $n\rightarrow\infty$.
  3. (3)
    This follows from the above proofs of (1) and (2).
In view of Theorem 5 and Lemma 6, the results of [KNS] can be revisited and recovered within our framework. As this would amount to a repetition of known results, we present the derivation primarily for conceptual clarity. The symmetric structure of our approach naturally lends itself to generalization, allowing for a direct extension to unitary transformations beyond the Fourier transform.
2. Uniqueness Pairs for Unitary Transformations beyond the Fourier Transform
We generalize Theorem 5 and provide a blueprint for identifying uniqueness pairs associated with a broader class of unitary transformations beyond the Fourier transform. Our derivation is based on the following assumptions:
  1. (1)
    There exists a subspace $\mathcal{X}\subset L^{2}(\mathbb{R})$ that arises as the domain and image of a densely-defined, essentially self-adjoint and positive operator $X$. The unitary transformation $U$ acts on this space, $U:\mathcal{X}\rightarrow\mathcal{X}$.
  2. (2)
    The operator $X$ is of the form [^2] Since we apply unitary transformations on various domains, we will not distinguish variables using the notation $x$ and $k$ as in the case of the Fourier transform. Henceforth, we will often use the same variable name on both sides.
    $X=\frac{1}{2}(M_{x^{2}}+U^{*}M_{x^{2}}U),$
    where $M_{x^{2}}$ is the multiplication operator $f\mapsto M_{x^{2}}(f)=x^{2}f$ on $\mathcal{X}$.
The assumptions are consistent. If a function $m:\mathbb{R}\rightarrow\mathbb{R}$ is measurable, then the multiplication operator $f\mapsto M_{m}(f)=mf$ with domain $\mathcal{D}(M_{m})=\{f\in L^{2}(\mathbb{R}):mf\in L^{2}(\mathbb{R})\}$ is self-adjoint on $L^{2}(\mathbb{R})$. If $\tilde{U}$ is a unitary transformation of $L^{2}(\mathbb{R})$ then $\tilde{U}^{*}M_{m}\tilde{U}$ is self-adjoint with domain $\tilde{U}^{*}\mathcal{D}(M_{m})$. By the Kato-Rellich theorem, $X$ is also self-adjoint with suitable domain in $L^{2}(\mathbb{R})$. As the average of two positive self-adjoint operators, $X$ is positive.
In analogy to the harmonic oscillator with infinite potential barriers, we consider the confined operators
$X_{[a,b]}=\begin{cases}X\ \textnormal{if}\ x\in[a,b],\\ \infty\ \textnormal{else},\end{cases}\ \textnormal{and}\quad(UXU^{*})_{[a,b]}=\begin{cases}UXU^{*}\ \textnormal{if}\ x\in[a,b],\\ \infty\ \textnormal{else}.\end{cases}$
Our discussion relies on spectral estimates to obtain lower bounds of the Raleigh coefficients
$\frac{\int_{a}^{b}\bar{f}X_{[a,b]}f\textnormal{d}x}{\int_{a}^{b}|f|^{2}\textnormal{d}x}\geq E_{[a,b]}^{X,(0)}\quad\textnormal{and}\quad\frac{\int_{a}^{b}\bar{f}(UXU^{*})_{[a,b]}f\textnormal{d}x}{\int_{a}^{b}|f|^{2}\textnormal{d}x}\geq E_{[a,b]}^{UXU^{*},(0)},\ f\neq 0.$
An important consideration in this context is the spectral nature of the involved operators. While $H$ and $H_{[a,b]}$ have purely discrete spectra, this need not hold for $X$ and $X_{[a,b]}$. However, if $X_{[a,b]}$ and $(UXU^{*})_{[a,b]}$ do have discrete spectra, the estimates become eigenvalue bounds, where $E^{X,(0)}_{[a,b]}$ and $E^{UXU^{*},(0)}_{[a,b]}$ denote the lowest eigenvalues. For completeness, let us mention a few important situations, where spectra are discrete.
Remark 7.
If $X$ has purely discrete spectrum then $X_{[a,b]}$ and $(UXU^{*})_{[a,b]}$ have purely discrete spectra.
Proof.
It is a standard fact that a self-adjoint operator on a Hilbert space has discrete spectrum if and only if it has compact resolvent. Using the resolvent identity one verifies that $U^{*}M_{x^{2}}U-M_{x^{2}}$ is relatively compact with respect to $M_{x^{2}}$. Let $P$ denote the projector onto $L^{2}([a,b])$, $X_{[a,b]}=PXP$. The operator $PM_{x^{2}}P$ is bounded and the operator $PU^{*}M_{x^{2}}UP$ is compact as a consequence of the mentioned relative compactness. Thus $X_{[a,b]}$ is the sum of a bounded and a compact operator and thus has a compact resolvent. The same reasoning applies to $(UXU^{*})_{[a,b]}$. ∎
Remark 8 ([RS1]).
Let $Y=-\frac{\partial^{2}}{\partial x^{2}}+V(x)$ be a Schrödinger operator on $L^{2}(\mathbb{R})$, where the potential $V:\mathbb{R}\rightarrow\mathbb{R}\cup\{\infty\}$ is measurable and bounded from below. If $V(x)\rightarrow\infty$ as $|x|\rightarrow\infty$ or $V(x)=\infty$ for $x\notin(a,b)$ then $Y$ has purely discrete spectrum.
Thus, if $X$ and $UXU^{*}$ are Schrödinger operators of the above type, the respective confined operators will automatically have discrete spectra. There is a variety of extensions of this result, most notably for second-order uniformly elliptic operators with Dirichlet boundary conditions on compact domains [D].
In what follows we will also consider subspaces $\mathcal{X}$ that arise via the embedding of functions defined on a subset $S\subset\mathbb{R}$ into $L^{2}(\mathbb{R})$. For instance, the Hankel transform (see below) is unitary on $L^{2}(\mathbb{R}^{+})$, which we interpret as a subspace of $L^{2}(\mathbb{R})$ via extension by zero.
Definition 9.
Let $(\Lambda,M)$ be a pair of countable subsets of $S\subset\mathbb{R}$ enumerated in increasing order with elements $(\lambda_{j})_{j\in\mathbb{Z}}$ and $(\mu_{j})_{j\in\mathbb{Z}}$, respectively. The pair $(\Lambda,M)$ has uniformly supercritical $X$-ground state energy for the transformation $U$ if
$\inf_{j\in\mathbb{Z}}\frac{\sqrt{E_{[\lambda_{j},\lambda_{j+1}]}^{X,(0)}}}{\max\{|\lambda_{j}|,|\lambda_{j+1}|\}}>1\ \textnormal{and}\ \inf_{j\in\mathbb{Z}}\frac{\sqrt{E_{[\mu_{j},\mu_{j+1}]}^{UXU^{*},(0)}}}{\max\{|\mu_{j}|,|\mu_{j+1}|\}}>1.$
We define the notions of super- and subcritical $X$-ground state energy, mutatis mutandis; compare also Definition 2.
Theorem 10.
Let the pair $(\Lambda,M)$ have uniformly supercritical $X$-ground state energy. Then the only function $f\in\mathcal{X}$ such that $f(\lambda_{i})=U[f](\mu_{i})=0$ for all $i\in\mathbb{Z}$ is $f=0$.
Proof.
We follow the proof of Theorem 5. Endow $\mathcal{X}$ with the norm $f|\!|_{\mathcal{X}}^{2}=<f|Xf>$. Then
[…]
$\geq\sum_{j\in\mathbb{Z}}\left(E_{[\lambda_{j},\lambda_{j+1}]}^{X,(0)}\int_{\lambda_{j}}^{\lambda_{j+1}}|f|^{2}\textnormal{d}x+E_{[\mu_{j},\mu_{j+1}]}^{UXU^{*},(0)}\int_{\mu_{j}}^{\mu_{j+1}}|U({f})|^{2}\textnormal{d}x\right)$
$\geq\sum_{j\in\mathbb{Z}}\left(\frac{E_{[\lambda_{j},\lambda_{j+1}]}^{X,(0)}}{\max\{|\lambda_{j}|^{2},|\lambda_{j+1}|^{2}\}}\int_{\lambda_{j}}^{\lambda_{j+1}}x^{2}|f|^{2}\textnormal{d}x+\frac{E_{[\mu_{j},\mu_{j+1}]}^{UXU^{*},(0)}}{\max\{|\mu_{j}|^{2},|\mu_{j+1}|^{2}\}}\int_{\mu_{j}}^{\mu_{j+1}}\overline{U(f)}x^{2}U(f)\textnormal{d}x\right)$
Akin to assuming uniformly supercritical $X$-ground state energy of $(\Lambda,M)$, we assume that the quantities in Definition 9 are bounded from below by $l$. It follows that
[…]
If $l>1$ this is a contradiction unless $f=0$. ∎
2.1. The Fractional Fourier Transform
The fractional Fourier transform $F_{\alpha}$ generalizes the Fourier transform by providing a representation that corresponds to an intermediate rotation by an angle $\alpha\in[0,2\pi)$ in the time-frequency domain. Numerous works are devoted to the study of the fractional Fourier transform and its applications; see, e.g., [YZ, OZK] for a review and references, and [J] for a discussion of some underlying technical principles.
On the Fourier-symmetric Sobolev space $\mathcal{H}$ the fractional Fourier transform can be defined by the relations
${F_{\alpha}}[f](k)=e^{-i\alpha H}[f](k),\quad f(x)=e^{i\alpha H}[F_{\alpha}f](x).$
If $\alpha=0$ no rotation occurs and we obtain the identity map, if $\alpha=\frac{\pi}{2}$ we recover the standard Fourier transform and $\alpha=\pi$ corresponds to spatial reflection: $F_{\pi/2}[f](x)\sim F[f](x)$, $F_{\pi}[f](x)\sim f(-x)$, both up to insignificant phase factors. The Hermite functions $\phi_{n}$ are eigenvectors of the fractional Fourier transform with eigenvalues $e^{-i\alpha(n+1/2)}$ and $X=\frac{1}{2}(M_{x^{2}}+F_{\alpha}^{*}M_{x^{2}}F_{\alpha})$ is essentially self-adjoint on $\mathcal{H}$. Writing $p=i\frac{\partial}{\partial x}$, $M_{x,\alpha}=F^{*}_{\alpha}M_{x}F_{\alpha}$ and $p_{\alpha}=F^{*}_{\alpha}pF_{\alpha}$, we recall the standard commutator relations $M_{x},p]=i$, $H,M_{x}]=ip$ and $H,p]=-iM_{x}$ and the Baker-Campbell-Hausdorff formulas
$\frac{\partial}{\partial\alpha}M_{x,\alpha}=i[H,M_{x,\alpha}]=p_{\alpha},$
$\frac{\partial}{\partial\alpha}p_{\alpha}=i[H,p_{\alpha}]=-M_{x,\alpha}.$
Solving this system yields
$M_{x,\alpha}$ $=\cos(\alpha)M_{x}+\sin(\alpha)p,$
$p_{\alpha}$ $=-\sin(\alpha)x+\cos(\alpha)p,$
which justifies that the fractional Fourier transform rotates the phase space and provides a quadratic-form representation for $X$,
$2X=(1+\cos^{2}(\alpha))M_{x^{2}}+\sin^{2}(\alpha)p^{2}+\sin(\alpha)\cos(\alpha)(M_{x}p+pM_{x}).$
It is a standard fact in the theory of oscillatory quantum systems that if a quadratic form in $(M_{x},p)$ is positive definite, then it can be transformed to a multiple of the Hamiltonian of the standard harmonic oscillator using a symplectic transformation $S:(M_{x},p)\mapsto(M_{x^{\prime}},p^{\prime})$, $M_{x},p]=[M_{x^{\prime}},p^{\prime}]$, see [GF, Chapter 4] and Williamson’s theorem [JW]. This implies
$2X=|\sin(\alpha)|(M_{x^{\prime}}^{2}+p^{\prime 2})$
with eigenvalues
$E_{n}^{X}=E_{n}^{F_{\alpha}XF_{\alpha}^{*}}=|\sin(\alpha)|\left(n+\frac{1}{2}\right).$
Thus, up to a scaling factor, the fractional Fourier transform is equivalent to the classical Fourier transform for the study of uniqueness pairs. This leads to a characterization of uniqueness pairs for the fractional Fourier transform following the arguments of Section 1 and [KNS, Theorem 1].
Corollary 11 (Uniqueness Pairs of the Fractional Fourier Transform).
If one of the following conditions holds then $(\Lambda,M)$ is a uniqueness pair for the fractional Fourier transform: the only function $f\in\mathcal{H}$ such that $f(\lambda_{i})=F_{\alpha}[f](\mu_{i})=0$ for all $i\in\mathbb{Z}$ is $f=0$.
  1. (1)
    $(\Lambda,M)$ has uniformly supercritical $X$-ground state energy
    $\inf_{j\in\mathbb{Z}}\frac{\sqrt{E_{[\lambda_{j},\lambda_{j+1}]}^{(0)}}}{\max\{|\lambda_{j}|,|\lambda_{j+1}|\}}>\frac{1}{|\sin(\alpha)|}\ \textnormal{and}\ \inf_{j\in\mathbb{Z}}\frac{\sqrt{E_{[\mu_{j},\mu_{j+1}]}^{(0)}}}{\max\{|\mu_{j}|,|\mu_{j+1}|\}}>\frac{1}{|\sin(\alpha)|},$
    where $E_{[a,b]}^{(0)}$ stands for the ground state energy of the Schrödinger equation (1.1).
  2. (2)
    The pair $(\Lambda,M)$ has supercritical $X$-ground state energy
    $\lim\sup|\lambda_{j}|(\lambda_{j+1}-\lambda_{j})<\pi|\sin(\alpha)|\ \textnormal{and}\ \lim\sup|\mu_{j}|(\mu_{j+1}-\mu_{j})<\pi|\sin(\alpha)|.$
On the other hand, if the pair $(\Lambda,M)$ has subcritical $X$-ground state energy,
$\lim\inf|\lambda_{j}|(\lambda_{j+1}-\lambda_{j})>\pi|\sin(\alpha)|\ \textnormal{and}\ \lim\inf|\mu_{j}|(\mu_{j+1}-\mu_{j})>\pi|\sin(\alpha)|,$
then it is not a uniqueness pair for the fractional Fourier transform.
It is interesting that Corollary 11 interpolates between the cases of minimal sufficient density for a uniqueness pair in the case of the standard Fourier transform to the requirement of infinite density when $\alpha$ approaches 0 or $\pi$:
$\lim\sup|\lambda_{j}|(\lambda_{j+1}-\lambda_{j})=0\ \textnormal{and}\ \lim\sup|\mu_{j}|(\mu_{j+1}-\mu_{j})=0.$
2.2. The Hankel transform
The Hankel transform of order $\nu>-\frac{1}{2}$, in the unitary convention, is defined as
$H_{\nu}[f](k)=\int_{\mathbb{R}^{+}}f(r)J_{\nu}(kr)\sqrt{kr}\textnormal{d}r,\ \textnormal{and}\ f(r)=H_{\nu}[H_{\nu}[f]](r),\quad f\in L^{2}(\mathbb{R}+),$
where $J_{\nu}$ denotes the Bessel function of first kind of order $\nu$. $H_{\nu}$ is unitary and involutory. We consider formally on $L^{2}(\mathbb{R}^{+})$ the “radial Hamiltonian” [^3] This operator is related to the Hamiltonian of the radial part of the wave function of the electron in a Hydrogen atom [NS].
$X=D_{\nu}=\frac{1}{2}\left(M_{r^{2}}+H_{\nu}M_{r^{2}}H_{\nu}\right),$
which is essentially self-adjoint and positive on a domain $\mathcal{D}\subset L^{2}(\mathbb{R}^{+})$. Endow $\mathcal{D}$ with the norm $f|\!|_{\mathcal{D}}^{2}=<f|D_{\nu}f>$.
Proposition 12.
The following assertions hold for the Hankel-symmetric space $\mathcal{D}$:
  1. (1)
    $\mathcal{D}\subset L^{2}(\mathbb{R}^{+})$ and $\mathcal{D}$ is the domain and image of the essentially self-adjoint and positive operator
    $D_{\nu}=\frac{1}{2}(L_{\nu}+M_{r^{2}})\quad\textnormal{with}\quad L_{\nu}=-\left(\frac{\partial^{2}}{\partial r^{2}}-\frac{\nu^{2}-\frac{1}{4}}{r^{2}}\right).$
  2. (2)
    $D_{\nu}$ has eigenvectors that are Laguerre functions of the form
    $\varphi_{n,\nu}=r^{\nu+\frac{1}{2}}e^{-\frac{r^{2}}{2}}L_{n,\nu}(r^{2}),\ n\geq 0,$
    where $L_{n,\nu}$ denote the associated Laguerre polynomials. The corresponding eigenvalues are $E_{n,\nu}=(2n+\nu+1)$.
  3. (3)
    The Laguerre functions form an orthogonal basis for $\mathcal{D}$.
  4. (4)
    $D_{\nu}$ is self-dual under the Hankel transform: ${H_{\nu}}D_{\nu}{H_{\nu}}^{-1}=D_{\nu}$.
  5. (5)
[…]
These facts are known but as notational conventions vary, we provide short proofs of (1), (2), (3) for the reader’s convenience. (4) and (5) are obvious. Refer to [GW] for details.
Proof.
Recall that the Bessel functions of the first kind solve the Bessel differential equation
$\frac{\partial^{2}J_{\nu}}{\partial r^{2}}+\frac{1}{r}\frac{\partial J_{\nu}}{\partial r}+\left(1-\frac{\nu^{2}}{r^{2}}\right)J_{\nu}=0,\quad x>0.$
  1. (1)
    Substituting $x=rk$ and applying the product rule, Bessel’s equation transforms to an eigenvalue equation
    $\left[-\frac{\partial^{2}}{\partial r^{2}}+\frac{\nu^{2}-\frac{1}{4}}{r^{2}}\right]\left(\sqrt{kr}J_{\nu}(kr)\right)=k^{2}\left(\sqrt{kr}J_{\nu}(kr)\right).$
    Applying the Hankel transform to this identity yields
    $H_{\nu}[r^{2}f](k)=-\left(\frac{\partial^{2}}{\partial k^{2}}-\frac{\nu^{2}-\frac{1}{4}}{k^{2}}\right)H_{\nu}[f](k).$
    Thus $H_{\nu}M_{r^{2}}H_{\nu}=L_{\nu}$ as claimed.
  2. (2)
    With ladder operators chosen as
    $a_{\nu}=\frac{\partial}{\partial r}+r-\frac{\nu+1/2}{r},\quad a_{\nu}^{\dagger}=-\frac{\partial}{\partial r}+r-\frac{\nu+1/2}{r},$
    direct computation shows $D_{\nu}=\frac{1}{2}\left(a_{\nu}^{\dagger}a_{\nu}+2\nu+2\right)$. Making use of the Laguerre identities
    $\frac{\partial}{\partial x}L_{n,\nu}(x)$ $=-L_{n-1,\nu+1}(x),$
    $L_{n,\nu}(x)$ $=L_{n,\nu+1}(x)-L_{n-1,\nu+1}(x),$
    $nL_{n,\nu}(x)$ $=(n+\nu)L_{n-1,\nu}(x)-xL_{n-1,\nu+1},$
    one verifies the action of ladder operators
    $a_{\nu}\varphi_{n,\nu}=-2\varphi_{n-1,\nu+1},\quad a^{\dagger}_{\nu}\varphi_{n-1,\nu+1}=-2n\varphi_{n,\nu}.$
    This yields the claimed eigenvector-eigenvalue relationship for $n,\nu$.
  3. (3)
    This is a consequence of the orthogonality of the associated Laguerre polynomials $L_{n,\nu}(x)$ with respect to the weighting function $x^{\nu}e^{-x}$.
We can estimate the energy levels $E^{D_{\nu},(n)}_{[a,b]}$ of the confined radial Hamiltonian $D_{\nu,[a,b]}$ on $a,b]\subset{\mathbb{R}^{+}}$ following the reasoning of Section 1,
$D_{\nu,[a,b]}^{down}=\begin{cases}-\frac{1}{2}\frac{\partial^{2}}{\partial r^{2}}+\frac{1}{2}\min\left\{\frac{\nu^{2}-\frac{1}{4}}{r^{2}}+r^{2}\right\}\ \textnormal{if}\ r\in[a,b],\\ \infty\ \textnormal{else},\end{cases}\quad\textnormal{and}\quad D_{\nu,[a,b]}>D_{n,[a,b]}^{down},$
$D_{\nu,[a,b]}^{up}=\begin{cases}-\frac{1}{2}\frac{\partial^{2}}{\partial r^{2}}+\frac{1}{2}\max\left\{\frac{\nu^{2}-\frac{1}{4}}{r^{2}}+r^{2}\right\}\ \textnormal{if}\ r\in[a,b],\\ \infty\ \textnormal{else},\end{cases}\quad\textnormal{and}\quad D_{\nu,[a,b]}<D_{\nu,[a,b]}^{up}.$
For the ground state energy of the confined radial harmonic oscillator this implies
$\frac{\pi^{2}}{2(b-a)^{2}}+\frac{1}{2}\max\left\{\frac{\nu^{2}-\frac{1}{4}}{r^{2}}+r^{2}\right\}>E^{D_{\nu},(0)}_{[a,b]}>\frac{\pi^{2}}{2(b-a)^{2}}+\frac{1}{2}\min\left\{\frac{\nu^{2}-\frac{1}{4}}{r^{2}}+r^{2}\right\}.$
As compared to the confined Hamiltonian of Section 1 the ground state energy level is shifted upwards. Thus, the $D_{\nu}$-ground energy can be bounded from below
$E^{D_{\nu},(0)}_{[a,b]}\;\geq\;\frac{\pi^{2}}{2(b-a)^{2}}+\frac{1}{2}\min\!\left\{\frac{\nu^{2}-\tfrac{1}{4}}{a^{2}}+a^{2},\;\frac{\nu^{2}-\tfrac{1}{4}}{b^{2}}+b^{2}\right\}\;\geq\;\sqrt{\nu^{2}-\tfrac{1}{4}}.$
This leads to a characterization of uniqueness pairs for the Hankel transform by Theorem 10.
Corollary 13 (Uniqueness Pairs of the Hankel Transform).
If the pair $(\Lambda,M)$ has uniformly supercritical $D_{\nu}$-ground state energy, then it is a uniqueness pair for the Hankel transform: the only function $f\in\mathcal{D}$ such that $f(\lambda_{i})=H_{\nu}[f](\nu_{i})=0$ for all $i\in\mathbb{Z}$ is $f=0$. If
$\lim\inf\frac{1}{\lambda_{j+1}^{2}}\left(\frac{\pi^{2}}{2(\lambda_{j+1}-\lambda_{j})^{2}}+\frac{1}{2}\min\!\left\{\frac{\nu^{2}-\tfrac{1}{4}}{\lambda_{j}^{2}}+\lambda_{j}^{2},\;\frac{\nu^{2}-\tfrac{1}{4}}{\lambda_{j+1}^{2}}+\lambda_{j+1}^{2}\right\}\right)$ $>1,$
$\lim\inf\frac{1}{\mu_{j+1}^{2}}\left(\frac{\pi^{2}}{2(\mu_{j+1}-\mu_{j})^{2}}+\frac{1}{2}\min\!\left\{\frac{\nu^{2}-\tfrac{1}{4}}{\mu_{j}^{2}}+\mu_{j}^{2},\;\frac{\nu^{2}-\tfrac{1}{4}}{\mu_{j+1}^{2}}+\mu_{j+1}^{2}\right\}\right)$ $>1,$
then the pair $(\Lambda,M)$ has uniformly supercritical $D_{\nu}$-ground state energy.
Along the lines of Corollary 6 it holds that if $(\Lambda,M)$ is uniformly supercritical, then $(\Lambda,M)$ has uniformly supercritical $D_{\nu}$-ground state energy with the exception of maximally a finite number of intervals.
3. Conclusion
In summary, this work develops a spectral framework for the analysis of uniqueness pairs associated with unitary transforms arising in signal processing. The proposed approach provides a unified treatment of two-sided sampling and related recovery problems. Our results show that uniqueness pairs admit a natural and tractable formulation within a general spectral setting. The flexibility and scope of the framework are illustrated through applications to the fractional Fourier and Hankel transforms.
Acknowledgements
This work was carried out during a research stay hosted by Andrii Bondarenko and Kristian Seip. The author is grateful to them for valuable discussions, and also wishes to thank the Department of Mathematical Sciences at NTNU for their warm hospitality and supportive environment. This work was supported by the Swiss National Science foundation, SNF grant No. CRSK-2_229036.